OpenStudy (kewlgeek555):

9x + 18 = 28x – 4 Steps Please.

OpenStudy (mathstudent55):

You need all terms with x on one side and all numbers on the other.

OpenStudy (kewlgeek555):

Okay...so, by doing that. What do I have to do? I mean Should I combine like terms like 28x - 4 = 24 x and 9 + 18 = 27x, which makes my equation 24x = 27x?

OpenStudy (anonymous):

9x+28x=23x 18-4=14 23x/23=0 23/14=1 x=1

OpenStudy (mathstudent55):

Like terms are terms that have the same variable part. On each side, the terms are not like terms and can't be combined.

OpenStudy (mathstudent55):

9x + 18 = 28x – 4 9x and 18 are not like terms and can't be combined. Likewise, 28x and -4 are not like terms and can't be combined.

OpenStudy (mathstudent55):

Since we want all x-terms on one side, let's move all x-terms to the left side. Since 28x is being added to -4 on the right side, we subtract 28x from both sides. 9x + 18 - 28x = 28x - 4 - 28x 9x - 28x + 18 = 28x - 28x - 4 Now we have like terms on each side that can be combined together, so we do it. -19x + 18 = -4 Now we onlt have x on the left side. We want the number terms on the right side. Since 18 is being added to -19, we subtract 18 from both sides.

OpenStudy (kewlgeek555):

Okay, so I got what you meant and this is my work and what I got: 9x-28x=-4-18 -19x=-22 x=1.15789473684

OpenStudy (mathstudent55):

-19x + 18 - 18 = -4 - 18 We combine like terms on both sides: -19x = -22 Now we have all x-terms on the left side and all number terms on the right side. x is being multiplied by -19, so to undo that, we divide both sides by -19: \(\dfrac{-19x}{-19} = \dfrac{-22}{-19} \) We reduce the fractions on both soides to get: \(x = \dfrac{22}{19} \) This is the final answer.

OpenStudy (kewlgeek555):

Okay, so we got the right answer. Yay. Thanks.

OpenStudy (mathstudent55):

wlcm